9. Let Y be a random variable who probability density function is given by P(Y- y) = 옮 when y is an intger between 1 and n inclusive. What should n be for this to be valid probability density function? Compute E(Y), V(Y), E(17Y - T) and 10, 10% of bottles produced at a factory have cracks. If two bottles are selected, find the mean and variance of the number of cracked bottles selected.
Answer #1

9) here as sum of probabilities of all sample space =1

hence (1/10+2/10+..+n/10)=1

n*(n+1)/20 =1

n*(n+1)=20

from above n=4

x f(x) xP(x) x2P(x)
1 0.100 0.100 0.100
2 0.200 0.400 0.800
3 0.300 0.900 2.700
4 0.400 1.600 6.400
total 3.000 10.000
E(x) =μ= ΣxP(x) = 3.0000
E(x2) = Σx2P(x) = 10.0000
Var(x)=σ2 = E(x2)-(E(x))2= 1.0000

E(Y)=3

Var(Y)=1.00

E(17Y-\pi)=17*E(Y)-\pi=17*3-\pi =51-\pi

E(1/Y)=(1/1)*(1/10)+(1/2)*(2/10)+(1/3)*(3/10)+(1/4)*(4/10)=0.4

10)

here this is binomial distribution with parameter n=2 and p=0.1

mean =np=2*0.1=0.2

variance =np(1-p)=2*0.1*(1-0.1)=0.18